# Situationally-Aware Dynamics Learning

### v2

@@ II-C1 Model-Free Reinforcement Learning:
II-C2 Model-Based Reinforcement Learning:
[havens2019learning] recognize that model-based methods are often restricted to tasks with limited dimensions or rely on certain assumptions about the system’s dynamics. Furthermore, they argue that high-dimensional models are challenging for planning due to increased compounding error, as also argued in [stachowicz2024racer, talvitie2014model, talvitie2017self]. Thus, they compress the state space to avoid learning unimportant features that could only contribute to compounding error during planning. Specifically, they propose to learn a state space encoder, a dynamics model in the latent space and action space, as well as a reward predictor. They place special emphasis in optimizing for reward prediction performance to encourage better compression of the state space. Closer to our work, [sharma2019dynamics] combine MBRL with primitives learned via MFRL to make model-based planning easier. In particular, they simultaneously identify “predictable behaviors” and the corresponding dynamics. Their key idea is to learn skills that are optimized to be easy to identify and for which a simpler accompanying dynamics model can be obtained. Finally, during deployment, the skill and corresponding dynamics model become a set of primitives that can be composed to solve the commanded task via model-based planning. In [sharma2020emergent], it is shown that this framework can be leveraged for RL in the real world. [lee2020context] introduces Context-aware Dynamics Models (CaDM) to generalize across different dynamics. They learn a context encoder, forward and backward dynamics models. Both dynamics models conditioned by the latent vector in charge of encoding dynamics-specific information.
@@ II-C3 Situational Awareness:
Other methods that do not rely on RL but leverage representations of the system’s situation include pervasive computing, where situation identification is used to abstract noisy sensor data into high-level concepts [ye2012situation]. For them, each situation is defined according to how the data is semantically interpreted by the users, which could require preliminary domain knowledge to obtain them. On the other hand, our work automatically builds situations by modelling the underlying data generation process that best explains the present sensor measurements. langari2005intelligent use situation identification to determine a policy for intelligent energy management in parallel hybrid vehicles. They do it by abstracting environment, driving style and car operating mode information using short term statistical features. However, they learn to classify the situations in a supervised manner and selecting which statistics to use requires task-specific expert-knowledge, which we avoid by modeling the UDGP of the transition dynamics. Other robot applications involve human-robot-cooperation to determine whether the robot maintains SA based on surprise [dahn2018situation].
II-D Generalized Hidden Parameter Markov Decision Process (GHP-MDP)
@@
[perez2020generalized] introduce GHP-MDPs and develop a learning-based method where agents can generalize, in simulation, to the same family of tasks using transition models whose input is augmented with estimated latent factors. They model the latent factors with a diagonal Gaussian while minimizing the Kullback–Leibler divergence between the variational and prior distributions. The main difference to our approach, is that we detect the latent representation online, which allows us to directly augment the (state, action, next-state)-tuples for fine-tuning the dynamics with the symbol assigned to each representation of the transition distribution. Furthermore, we use a full-rank covariance matrix, instead of a less expressive diagonal covariance.
@@
Related work also includes Hidden Parameter MDPs (HiP-MDPs) which assume the agent knows what it does not know. Specifically, it does not know the value of some critical environment variables and focuses on estimating these [doshi2016hidden]. [zhang2020learning] learn state abstractions to enable multi-task RL. However, they assume the latent parameter is fixed for the duration of the episode, while in our case we are interested in detecting and accommodating to changes in the latent parameters during the episode. [killian2017robust] extend this framework for robust and efficient transfer learning to subtle variations among tasks using Bayesian Neural Networks. [costen2023planning] develop a method to express the transition probabilities in closed-form as a polynomial of the latent parameters, as well as maintaining a closed-form belief distribution of the latent parameters. They present promising results in two simulation tasks. In our case, the transition model’s input is augmented with the symbol of the detected latent representation, since we do not assume prior-knowledge about the variables that the agent will require to obtain a better model of the dynamics at deployment time.
II-E Unstructured Terrain Navigation
@@
The validation of our approach is through the challenging task of unstructured terrain navigation, where unmodeled and unmeasured terrain characteristics can significantly impact the robot’s motion. Unstructured terrain navigation has been tackled mostly from two perspectives: adaptation and path planning. Works addressing this from the adaptive perspective typically devise methods that allow the robot to recover from unexpected or hard scenarios. For example, xu2024kernel proposes a method to bypass the precise modeling of the state transition function of an MDP with continuous state space and shows that it can be applied for off-road navigation and recovering from unexpected phenomena (e.g., wheel slippage). wang2021rough learns a probabilistic dynamics model trained to consider the propagating effects of uncertainty and uses it along a tracking controller and constrained optimization to find trajectories with low divergence. Terrain-aware methods have been recently developed to generate consistent navigation behaviors by learning a control offset from demonstrations [siva2019robot, siva2021enhancing]. On the other hand, the methods concerned with global planning typically use the robot’s sensors to analyze the terrain and try to find the trajectory that seems safer and easier to navigate [yoo2024traversability, jardali2024autonomous, leininger2024gaussian]. Other methods have done so by focusing on risk-awareness [sharma2023ramp, cai2022risk, cai2023probabilistic]. Learning-based approaches have considered ground-robot interaction and surface data [sathyamoorthy2022terrapn]. Our method can be framed at the intersection of both perspectives. In particular, we obtain dynamics models with greater robustness to terrain challenges and use those to make local motion plans that help the robot traverse safely under partial state information.
III Preliminaries
III-A Generalized Hidden Parameter Markov Decision Process (GHP-MDP)
@@
Markov Decision Processes (MDPs) provide a framework for modeling decision-making with uncertainty over the action outcomes due to an agent’s interaction with its environment [sutton2018reinforcement]. Hidden Parameter MDPs (HiP-MDP) extend MDPs to scenarios where predicting the future also requires a set of unobservable parameters $\Theta$ that correspond to the sufficient statistics to adapt to a new situation [doshi2016hidden]. In HiP-MDPs, $\Theta$ explicitly represent a critical variable of the environment; for example, physical properties like the terrain friction. This makes it necessary to know a-priori which missing variables need to be estimated by $\Theta$, which is a strong assumption we cannot fulfill when the robot is navigating unstructured terrains as we cannot anticipate every challenge it will face.
@@
[perez2020generalized] build upon HiP-MDPs and present GHP-MDPs to account for unobserved, latent parameters which influence the environment’s dynamics and reward function. These latent factors are not directly observable by the agent, but the agent must infer them through interaction with the environment to optimize its policy.
@@
A Generalized Hidden Parameter MDP (GHP-MDP) is defined by a tuple $\mathcal{M}_{\text{GHP}}=\left(S,A,\Theta,T,R,\gamma\right)$, where $S$ is the set of observable states and $A$ is the set of actions available to the agent. $\Theta$ is the set of latent variables that are not observable by the agent but influence the system’s dynamics and rewards. $T:S\times\Theta\times A\times S\to[0,1]$ is the transition function which gives the probability of transitioning from state $s$ to state $s^{\prime}$ under action $a$ and hidden parameter $\theta$. $R:S\times\Theta\times A\to\mathbb{R}$ is the reward function. And the discount factor $\gamma\in[0,1]$ modulates the agent’s desire for immediate and future rewards.
@@
As the agent in a GHP-MDP does not directly observe $\theta$, it must maintain a belief $b(\theta)$ over the possible values of the latent variables. In this work, we present a novel and efficient approach for representing $b(\theta)$ online and unsupervisedly as the distribution of $(s,a,s^{\prime})$-tuples being currently observed. Each distribution is then mapped to a symbol that will be used by $T$ to adapt to the current situation of the robot.
III-A1 Comparison to other MDP Models:
@@
Throughout the years multiple extensions of the MDP framework have been proposed. In the following, we highlight key differences to some of them to justify selecting a GHP-MDP to formulate the unstructured terrain navigation problem with unobservable and changing state variables:
@@
MDP [bellman1958dynamic]: The state $s$ is fully observable, and the environment’s dynamics are completely determined by the state and action, i.e., $T(s,a,s^{\prime})=P(s^{\prime}|s,a)$.
@@
POMDP [aastrom1965optimal, kaelbling1998planning]: The agent maintains a belief over the partially-observable state $s$. It must infer the state using the observations $o$. Conversely, in a GHP-MDP, we do not have uncertainty over the observable part of the state $s$, and the uncertainty arises from the unknown latent factors $\theta$ that are affecting the environment dynamics.
@@
Hidden Model MDP, Contextual MDP (cMDP) or Latent-Variable MDP (LMDP) [chades2012momdps, hallak2015contextual, kwon2021rl, steimle2021multi]: The model consists of a set of transition and reward functions from the same state and action space. The number of transition and reward functions is given by the number of contexts the agent could be in and the value function is the expectation over these, weighted according to a set of mixing weights. On the other hand, GHP-MDPs have a single transition and reward function whose input is augmented by the latent representation.
III-B Model-Based Reinforcement Learning
@@
An MBRL agent interacts with the environment by learning its model and using it to make decisions. A key difference with other system identification techniques is the exploratory nature of MBRL. Specifically, the agent acts on the environment while balancing exploration and exploitation [sutton2018reinforcement]. Every episode it visits more informative regions as it avoids what it expects to be known or irrelevant configurations of the state-action space, which translates to faster model learning via informed data collection. This approach contrasts with MFRL, where the agent directly learns a policy solely based on interactions without explicit knowledge of the environment’s dynamics.
@@
III-C Bayesian Online Changepoint Detection (BOCD)
The problem of changepoint detection is concerned with determining the point where the observed data distribution changes in an ordered set of measurements, such as in time-series [fearnhead2019changepoint]. BOCD was introduced by adams2007bayesian to tackle this problem by framing it as an estimation of the posterior distribution of the current “run length” $r_{t}$, meaning how likely it is that the measurement at time-step $t$ belongs to the same data generating process that started $r_{t}$ timesteps ago; while also obtaining the parameters $\boldsymbol{\eta}$ defining such process.
@@ III-C Bayesian Online Changepoint Detection (BOCD)
where $H(\tau)$ is the hazard function [forbes2011statistical]. Subsequently, finding the distribution of the run length $P(r_{t}|\boldsymbol{x}_{1:t})$, involves recursive computation over the run length and the observations:
@@
IV Method

### v3

@@ 2.3.1 Model-Free Reinforcement Learning:
2.3.2 Model-Based Reinforcement Learning:
Havens et al. (2019) recognize that model-based methods are often restricted to tasks with limited dimensions or rely on certain assumptions about the system’s dynamics. Furthermore, they argue that high-dimensional models are challenging for planning due to increased compounding error, as also argued in Stachowicz and Levine (2024); Talvitie (2014, 2017). Thus, they compress the state space to avoid learning unimportant features that could only contribute to compounding error during planning. Specifically, they propose to learn a state space encoder, a dynamics model in the latent space and action space, as well as a reward predictor. They place special emphasis in optimizing for reward prediction performance to encourage better compression of the state space. Closer to our work, Sharma et al. (2019) combine MBRL with primitives learned via MFRL to make model-based planning easier. In particular, they simultaneously identify “predictable behaviors” and the corresponding dynamics. Their key idea is to learn skills that are optimized to be easy to identify and for which a simpler accompanying dynamics model can be obtained. Finally, during deployment, the skill and corresponding dynamics model become a set of primitives that can be composed to solve the commanded task via model-based planning. In Sharma et al. (2020), it is shown that this framework can be leveraged for RL in the real world. Lee et al. (2020b) introduces Context-aware Dynamics Models (CaDM) to generalize across different dynamics. They learn a context encoder, forward and backward dynamics models. Both dynamics models conditioned by the latent vector in charge of encoding dynamics-specific information.
@@ 2.3.3 Situational Awareness:
Other methods that do not rely on RL but leverage representations of the system’s situation include pervasive computing, where situation identification is used to abstract noisy sensor data into high-level concepts Ye et al. (2012). For them, each situation is defined according to how the data is semantically interpreted by the users, which could require preliminary domain knowledge to obtain them. On the other hand, our work automatically builds situations by modelling the underlying data generation process that best explains the present sensor measurements. Langari and Won (2005) use situation identification to determine a policy for intelligent energy management in parallel hybrid vehicles. They do it by abstracting environment, driving style and car operating mode information using short term statistical features. However, they learn to classify the situations in a supervised manner and selecting which statistics to use requires task-specific expert-knowledge, which we avoid by modeling the UDGP of the transition dynamics. Other robot applications involve human-robot-cooperation to determine whether the robot maintains SA based on surprise Dahn et al. (2018).
2.4 Generalized Hidden Parameter Markov Decision Process (GHP-MDP)
@@
Perez et al. (2020) introduce GHP-MDPs and develop a learning-based method where agents can generalize, in simulation, to the same family of tasks using transition models whose input is augmented with estimated latent factors. They model the latent factors with a diagonal Gaussian while minimizing the Kullback–Leibler divergence between the variational and prior distributions. The main difference to our approach, is that we detect the latent representation online, which allows us to directly augment the (state, action, next-state)-tuples for fine-tuning the dynamics with the symbol assigned to each representation of the transition distribution. Furthermore, we use a full-rank covariance matrix, instead of a less expressive diagonal covariance.
@@
Related work also includes Hidden Parameter MDPs (HiP-MDPs) which assume the agent knows what it does not know. Specifically, it does not know the value of some critical environment variables and focuses on estimating these Doshi-Velez and Konidaris (2016). Zhang et al. (2020) learn state abstractions to enable multi-task RL. However, they assume the latent parameter is fixed for the duration of the episode, while in our case we are interested in detecting and accommodating to changes in the latent parameters during the episode. Killian et al. (2017) extend this framework for robust and efficient transfer learning to subtle variations among tasks using Bayesian Neural Networks. Costen et al. (2023) develop a method to express the transition probabilities in closed-form as a polynomial of the latent parameters, as well as maintaining a closed-form belief distribution of the latent parameters. They present promising results in two simulation tasks. In our case, the transition model’s input is augmented with the symbol of the detected latent representation, since we do not assume prior-knowledge about the variables that the agent will require to obtain a better model of the dynamics at deployment time.
2.5 Unstructured Terrain Navigation
@@
We validate our approach in the real world through the challenging task of unstructured terrain navigation, where unmodeled and unmeasured terrain characteristics can significantly impact the robot’s motion. Unstructured terrain navigation has been tackled mostly from two perspectives: adaptation and path planning. Works addressing this from the adaptive perspective typically devise methods that allow the robot to recover from unexpected or hard scenarios. For example, Xu et al. (2024a) proposes a method to bypass the precise modeling of the state transition function of an MDP with continuous state space and shows that it can be applied for off-road navigation and recovering from unexpected phenomena (e.g., wheel slippage). Wang et al. (2021) learns a probabilistic dynamics model trained to consider the propagating effects of uncertainty and uses it along a tracking controller and constrained optimization to find trajectories with low divergence. Terrain-aware methods have been recently developed to generate consistent navigation behaviors by learning a control offset from demonstrations Siva et al. (2019, 2021). Pokhrel et al. (2024) develop a multi-stage approach where terrain representations for offroad navigation are first learned in a self-supervised manner from visual, inertial and speed observation. Then in a subsequent stage, the terrain representation weights are frozen and the forward kinodinamic model is learned from a terrain interaction dataset. In contrast, our method learns the situation representations online and unsupervised. On the other hand, the methods concerned with global planning typically use the robot’s sensors to analyze the terrain and try to find the trajectory that seems safer and easier to navigate Yoo et al. (2024); Jardali et al. (2024); Leininger et al. (2024). For example, Castro et al. (2022) use self-supervised learning to combine exteroceptive and proprioceptive feedback to predict traversability costmaps. Other methods have done so by focusing on risk-awareness Sharma et al. (2023); Cai et al. (2022, 2023). Learning-based approaches have considered ground-robot interaction and surface data Sathyamoorthy et al. (2022). Our method can be framed at the intersection of both perspectives. In particular, we obtain dynamics models with greater robustness to terrain challenges and use those to make local motion plans that help the robot traverse safely under partial state information.
3 Preliminaries
3.1 Generalized Hidden Parameter Markov Decision Process (GHP-MDP)
@@
Markov Decision Processes (MDPs) provide a framework for modeling decision-making with uncertainty over the action outcomes due to an agent’s interaction with its environment Sutton and Barto (2018). Hidden Parameter MDPs (HiP-MDP) extend MDPs to scenarios where predicting the future also requires a set of unobservable parameters $\Theta$ that correspond to the sufficient statistics to adapt to a new situation Doshi-Velez and Konidaris (2016). In HiP-MDPs, $\Theta$ explicitly represent a critical variable of the environment; for example, physical properties like the terrain friction. This makes it necessary to know a-priori which missing variables need to be estimated by $\Theta$, which is a strong assumption we cannot fulfill when the robot is operating in unstructured environments as we cannot anticipate every challenge it will face.
@@
Perez et al. (2020) build upon HiP-MDPs and present GHP-MDPs to account for unobserved, latent parameters which influence the environment’s dynamics and reward function. These latent factors are not directly observable by the agent, but the agent must infer them through interaction with the environment to optimize its policy.
@@
A Generalized Hidden Parameter MDP (GHP-MDP) is defined by a tuple $\mathcal{M}_{\text{GHP}}=\left(S,A,\Theta,T,R,\gamma\right)$, where $S$ is the set of observable states and $A$ is the set of actions available to the agent. $\Theta$ is the set of latent variables that are not observable by the agent but influence the system’s dynamics and rewards. $T:S\times\Theta\times A\times S\to[0,1]$ is the transition function which gives the probability of transitioning from state $s\in S$ to state $s^{\prime}\in S$ under action $a\in A$ and hidden parameter $\theta$. $R:S\times\Theta\times A\to\mathbb{R}$ is the reward function. And the discount factor $\gamma\in[0,1]$ modulates the agent’s desire for immediate and future rewards.
@@
As the agent in a GHP-MDP does not directly observe $\theta$, it must maintain a belief $b(\theta)$ over the possible values of the latent variables. In this work, we present a novel, efficient and unsupervised approach for representing $b(\theta)$ online as the distribution of $(s,a,s^{\prime})$-tuples being currently observed. Each distribution is then mapped to a symbol that will be used by $T$ to adapt to the current situation of the robot.
3.1.1 Comparison to other MDP Models:
@@
Throughout the years multiple extensions of the MDP framework have been proposed. In the following, we highlight key differences to some of them to justify selecting a GHP-MDP to formulate the decision making problem in unstructured environments with unobservable and changing state variables:
@@
MDP Bellman (1958): The state $s$ is fully observable, and the environment’s dynamics are completely determined by the state and action, i.e., $T(s,a,s^{\prime})=P(s^{\prime}|s,a)$.
@@
POMDP Åström (1965); Kaelbling et al. (1998): The agent maintains a belief over the partially-observable state $s$. It must infer the state using the observations $o$. Conversely, in a GHP-MDP, we do not necessarily have uncertainty over the observable part of the state $s$, and the uncertainty arises from the unknown latent factors $\theta$ that are affecting the system dynamics.
@@
Hidden Model MDP, Contextual MDP (cMDP) or Latent-Variable MDP (LMDP) Chades et al. (2012); Hallak et al. (2015); Kwon et al. (2021); Steimle et al. (2021): The model consists of a set of transition and reward functions from the same state and action space. The number of transition and reward functions is given by the number of contexts the agent could be in and the value function is the expectation over these, weighted according to a set of mixing weights. On the other hand, GHP-MDPs have a single transition and reward function whose input is augmented by the latent factors representation.
3.2 Model-Based Reinforcement Learning
@@
An MBRL agent interacts with the environment by learning its model and using it to make decisions. A key difference with other system identification techniques is the exploratory nature of MBRL. Specifically, the agent acts on the environment while balancing exploration and exploitation Sutton and Barto (2018). Every episode it visits more informative regions as it avoids what it expects to be known or irrelevant configurations of the state-action space, which translates to faster model learning via informed data collection. This approach contrasts with MFRL, where the agent directly learns a policy solely based on interactions without explicit knowledge of the environment’s dynamics.
@@
3.3 Bayesian Online Changepoint Detection (BOCD)
The problem of changepoint detection is concerned with determining the point where the observed data distribution changes in an ordered set of measurements, such as in time-series Fearnhead and Rigaill (2019). BOCD was introduced by Adams and MacKay (2007) to tackle this problem by framing it as an estimation of the posterior distribution of the current “run length” $r_{t}$, meaning how likely it is that the measurement at time-step $t$ belongs to the same data generating process that started $r_{t}$ timesteps ago; while also obtaining the parameters $\boldsymbol{\eta}$ defining such process.
@@ 3.3 Bayesian Online Changepoint Detection (BOCD)
where $H(\tau)$ is the hazard function Forbes et al. (2011). Subsequently, finding the distribution of the run length $P(r_{t}|\boldsymbol{x}_{1:t})$, involves recursive computation over the run length and the observations:
@@
4 Method
We propose a method that learns a situationally-aware dynamics model by identifying latent, unobserved factors from online interaction data. We organize our presentation in a top-down manner. First, we describe the overall pipeline: a model-based reinforcement learning agent that leverages a learned representation of the current ‘situation’ to improve planning and control. This high-level overview explains how the system works as a whole. Second, we detail our core contribution: an online and unsupervised method for discovering these situations by detecting changes in the robot’s transition dynamics. This section provides the theoretical underpinnings of our approach.
4.1 Situationally-Aware Dynamics
For effective planning and control in changing or unstructured environments, an agent must account for hidden state variables that influence its dynamics. We address this by augmenting the standard MDP transition model with a parameter $\theta\in\Theta$ that represents these unmodeled world characteristics. This results in a GHP-MDP with transition function:
$T(s,\theta,a,s^{\prime})=P(S_{t+1}=s^{\prime}\mid S_{t}=s,\Theta_{t}=\theta,A_{t}=a),$ (8)
where $S_{t},\Theta_{t},A_{t}\text{ and }S_{t+1}$ are random variables representing the state of the world, the current representation of the latent factors, and the actions taken by the agent. The parameter $\theta$ encapsulates the current “situation” of the robot (e.g., traversing slippery ice versus firm ground in a UGV navigation scenario), allowing the dynamics model to adapt its predictions accordingly. An overview of our approach is shown in Figure 2.
In practical terms, we model the right-hand side of Eq. (8) as an ensemble $\mathcal{M}_{m}:S\times\Theta\times A\rightarrow S$ of $m$ probabilistic (Gaussian) neural networks. The ensemble’s input is augmented with a representation of the current situation, $\theta$, in addition to the states and actions. Together, these inputs enable the ensemble to learn the effects of the hidden state on the robot’s dynamics. This architecture choice estimates the predictive distribution over the next state and simultaneously considers the epistemic and aleatoric uncertainty Chua et al. (2018).
4.1.1 Training and Testing Setup:
The system operates in two distinct phases: training and testing. The training phase is designed to build the situation library $\boldsymbol{\Psi}$ and learn the dynamics model $\mathcal{M}_{m}$, and can be performed in either an active or passive manner.
In the active setting, the agent interacts directly with the environment. It learns by alternating exploration and exploitation, following the MBRL framework described in Algorithm 1. During this process, transition data is collected and our situation identification algorithm (detailed in Sec. 4.2) runs in parallel to discover new situations and populate the library $\boldsymbol{\Psi}$. The collected transitions are stored in a replay buffer $\mathcal{R}=\mathcal{R}\cup\{(s_{i},\theta_{i},a_{i},s_{i+1})\}_{i=1}^{N_{e}}$, where each tuple also includes the symbol $\theta_{i}$ representing the identified situation $\psi_{i}\in\boldsymbol{\Psi}$ and $N_{e}$ are the number of steps in the episode.
On the other hand, in the passive setting, the agent learns from a pre-existing, static dataset of transitions $\mathcal{D}=\{(s_{i},a_{i},s_{i+1})\}_{i=1}^{N_{d}}$. Here, no exploration occurs. Instead, to construct the situation-annotated replay buffer, we process the dataset $\mathcal{D}$ sequentially: as each transition is fed to our system, our online situation identification algorithm discovers the underlying situations present in the data and builds the library $\boldsymbol{\Psi}$. This allows us to annotate each tuple with its corresponding situation symbol $\theta_{i}$, creating the final replay buffer $\mathcal{R}=\{(s_{i},\theta_{i},a_{i},s_{i+1})\}_{i=1}^{N_{d}}$.
It is worth noting that a hybrid configuration is also possible: the agent can start from a pre-collected dataset (passive learning) and then transition to active data collection for further dynamics model refinement.
Regardless of the training paradigm, the resulting replay buffer $\mathcal{R}$ that will be used to learn the model should contain transitions annotated with situation symbols. Furthermore, given that the period of time the robot experiences in each situation is different, the replay buffer $\mathcal{R}$ will naturally have an unbalanced sample of situations. Thus, to ensure that the dynamics $\mathcal{M}_{m}$ are learned appropriately for all situations $\psi\in\boldsymbol{\Psi}$, we introduce a loss weighted by the cardinality of each situation. Specifically, we minimize the weighted Negative Log-Likelihood (NLL):
During the testing phase, the situation library $\boldsymbol{\Psi}$ is held fixed. The agent no longer creates new situations. Instead, it continuously identifies its current situation by comparing its local dynamics to the learned models in $\boldsymbol{\Psi}$ and selecting the best match (detailed in Sec. 4.2.3). This identified situation informs the dynamics model $\mathcal{M}_{m}$, allowing for adaptive control in response to changing conditions.
4.1.2 Planning and Control:
We use our situationally-aware dynamics model $\mathcal{M}_{m}$ together with the Model Predictive Path Integral (MPPI) method (Gandhi et al., 2021). The planning process involves generating a set of candidate action sequences over a time horizon $H$ and selecting the sequence that maximizes the expected return. The expected return is computed based on the predicted future states while assuming that the situation symbol $\theta$ remains constant throughout the rollouts. To select an action, MPPI optimizes the function:
$\mathcal{B}(\hat{s}_{t+i})=-\log\|\hat{s}_{t+i}-s_{obs}\|_{2}.$ (12)
Every timestep, the controller applies the first action in $a^{\star}_{t:t+H}$ and then re-evaluates and updates the action sequence, ensuring that the agent adapts to changes in the environment.
The effectiveness of the situationally-aware dynamics model described above hinges on its ability to accurately identify the current situation, $\psi$, and represent it as a symbol $\theta$. The following section presents our main theoretical contribution: a novel method for discovering and modeling these situations online and in an unsupervised fashion.
4.2 Online and Unsupervised Latent Factor Representation Learning
Figure 3: Motivating example showing the predictive benefit of conditioning on latent factors. Plotted are the marginal distributions of the change in state $\Delta s$ conditioned on each terrain type from a real-world UGV dataset; colored curves correspond to terrain-specific distributions and the dashed black curve shows the unconditioned (aggregated) estimate. The unconditioned estimate smooths over distinct, terrain-dependent modes and therefore fails to capture terrain-specific transition dynamics, illustrating why latent-factor conditioning yields more accurate predictions. See Appendix A for details.